# Solvable Tuple Patterns and Their Applications to Program Verification

### v1

@@ B.1 Extensions with Map
Let $f$ be $\{a\mapsto a,b\mapsto a\}$ and $M$ be:
$\left(\begin{array}[]{rrr}aa&\ aaaa\\ b&\ aa\\ \end{array}\right).$
We can infer the pattern $(x,{x}^{f}{x}^{f})$ as follows.
$((x_{1},x_{2}),[M/(x_{1},x_{2})])$
$\longrightarrow((x_{1},{x_{1}}^{f}x^{\prime}_{2}),[\left(\begin{array}[]{rrrr}aa&\ aa\\ b&\ a\\ \end{array}\right)/(x_{1},x^{\prime}_{2})])$ (R-MPrefix1)
$\longrightarrow((x_{1},{x_{1}}^{f}{x_{1}}^{f}x^{\prime\prime}_{2}),[\left(\begin{array}[]{rrrr}aa&\ \epsilon\\ b&\ \epsilon\\ \end{array}\right)/(x_{1},x^{\prime\prime}_{2})])$ (R-MPrefix1)
$\longrightarrow((x_{1},{x_{1}}^{f}{x_{1}}^{f}),[\left(\begin{array}[]{rrrr}aa\\ b\end{array}\right)/x_{1}])$ $\mbox{ (\hbox{\sc R-Epsilon})}.$
Remark B.1.
Without the rule R-MPrefix2 (and the corresponding rule PR-MPrefix2), the minimality would not hold. Let $f=\{a\mapsto c,b\mapsto d\}$ and $M$ be:
$\left(\begin{array}[]{rrr}cda&\ ab&\ c\end{array}\right).$
The pattern $(zx,y,z)$ is inferred as follows.
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])$
$\longrightarrow((x_{3}x^{\prime}_{1},x_{2},x_{3}),[\left(\begin{array}[]{rrrr}da&\ ab&\ c\end{array}\right)/(x_{1}^{\prime},x_{2},x_{3})])$ (R-Prefix)
$\;\mathbin{\ \not\!\!\longrightarrow}.$
It is not minimal as $({x}^{f}{y}^{f}x,xy,{x}^{f})$ can also be inferred:
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])$
$\longrightarrow(({x_{2}}^{f}x^{\prime}_{1},x_{2},x_{3}),[\left(\begin{array}[]{rrrr}a&\ ab&\ c\end{array}\right)/(x_{1}^{\prime},x_{2},x_{3})])$ (R-MPrefix1)
$\longrightarrow(({x_{2}}^{f}x^{\prime}_{1},x_{2},{x^{\prime}_{1}}^{f}x^{\prime}_{3}),[\left(\begin{array}[]{rrrr}a&\ ab&\ \epsilon\end{array}\right)/(x_{1}^{\prime},x_{2},x^{\prime}_{3})])$ (R-MPrefix1)
$\longrightarrow(({x^{\prime}_{1}}^{f}{x^{\prime}_{2}}^{f}x^{\prime}_{1},x^{\prime}_{1}x^{\prime}_{2},{x^{\prime}_{1}}^{f}x^{\prime}_{3}),[\left(\begin{array}[]{rrrr}a&\ b&\ \epsilon\end{array}\right)/(x_{1}^{\prime},x^{\prime}_{2},x^{\prime}_{3})])$ (R-Prefix)
$\longrightarrow(({x^{\prime}_{1}}^{f}{x^{\prime}_{2}}^{f}x^{\prime}_{1},x^{\prime}_{1}x^{\prime}_{2},{x^{\prime}_{1}}^{f}),[\left(\begin{array}[]{rrrr}a&\ b\end{array}\right)/(x_{1}^{\prime},x^{\prime}_{2})])$ (R-Epsilon)
$\;\mathbin{\ \not\!\!\longrightarrow}.$
In the presence of R-MPrefix2, the first reduction sequence can be extended as follows.
$((x_{1},x_{2},x_{3}),[M/(x_{1},x_{2},x_{3})])$
$\longrightarrow((x_{3}x^{\prime}_{1},x_{2},x_{3}),[\left(\begin{array}[]{rrrr}da&\ ab&\ c\end{array}\right)/(x_{1}^{\prime},x_{2},x_{3})])$ (R-Prefix)
$\longrightarrow(({x^{\prime}_{3}}^{f}x^{\prime}_{1},x^{\prime}_{3}x_{2}^{\prime},{x^{\prime}_{3}}^{f}),[\left(\begin{array}[]{rrrr}da&\ b&\ a\end{array}\right)/(x_{1}^{\prime},x^{\prime}_{2},x_{3}^{\prime})])$ (R-MPrefix2)
$\longrightarrow(({x^{\prime}_{3}}^{f}{x_{2}^{\prime}}^{f}x^{\prime\prime}_{1},x^{\prime}_{3}x_{2}^{\prime},{x^{\prime}_{3}}^{f}),[\left(\begin{array}[]{rrrr}a&\ b&\ a\end{array}\right)/(x_{1}^{\prime\prime},x^{\prime}_{2},x_{3}^{\prime})])$ (R-MPrefix1)
$\longrightarrow(({x^{\prime}_{3}}^{f}{x_{2}^{\prime}}^{f}x^{\prime\prime}_{1},x^{\prime}_{3}x_{2}^{\prime},{(x_{1}^{\prime\prime}x_{3}^{\prime\prime})}^{f}),[\left(\begin{array}[]{rrrr}a&\ b&\ \epsilon\end{array}\right)/(x_{1}^{\prime\prime},x^{\prime}_{2},x_{3}^{\prime\prime})])$ (R-Prefix)
$\longrightarrow(({x^{\prime}_{3}}^{f}{x_{2}^{\prime}}^{f}x^{\prime\prime}_{1},x^{\prime}_{3}x_{2}^{\prime},{x_{1}^{\prime\prime}}^{f}),[\left(\begin{array}[]{rrrr}a&\ b\end{array}\right)/(x_{1}^{\prime\prime},x^{\prime}_{2})])$ $\mbox{ (\hbox{\sc R-Epsilon})}.$
The failure of the minimality (without R-MPrefix2) can be attributed to that of the “weak confluence” property (Lemma A.5) needed for the proof of minimality. Consider $({x}^{f}{y}^{f}x,xy,{x}^{f})$. It is solvable without using PR-MPrefix2 since
$({x}^{f}{y}^{f}x,xy,{x}^{f})\leadsto(x,xy,{x}^{f})\leadsto(x,xy,\epsilon)\leadsto(x,y,\epsilon)\leadsto(x,y).$
However, the following reduction gets stuck (without using PR-MPrefix2) as follows.
$({x}^{f}{y}^{f}x,xy,{x}^{f})\leadsto({y}^{f}x,xy,{x}^{f})\mathbin{\ \not\!\leadsto}$
Thus, weak confluence of $\leadsto$ would fail without PR-MPrefix2. ∎

### v2

@@ Appendix B Verification of Functional Queues Using Piecewise Conjunctive Solvable Tuple Patterns
Let us consider the following OCaml program.
let initq = ([], [])
let rec reva l1 l2 =
  match l1 with
    [] -> l2
  | x::l1’ -> reva l1’ (x::l2)
let enq x (l1,l2) = (l1, x::l2)
let deq (l1,l2) =
  match l1 with
     (match reva l2 [] with
        [] -> ([], l1, l2)
      | x::l -> ([x], l, [])
  | x::l1’ -> ([x], l1’, l2)
let rec enqall l (l1,l2) =
  match l with
    [] -> (l1,l2)
  | x::l’ -> enqall l’ (enq x (l1,l2))
let rec deqall (l1,l2) =
  match deq(l1,l2) with
    ([], _, _) -> []
  | (l, l1’,l2’) -> l@(deqall (l1’, l2’))
let main l =
  assert(deqall (enqall l initq) = l)
Here, a queue is implemented as a pair of lists $(l_{1},l_{2})$ where $l_{1}l_{2}^{R}$ is the sequence of elements in the queue, so that the amortized cost of each enqueue or dequeue operation is $O(1)$ [48]. The dequeue function deq either returns [][\,] (when the queue is empty, i.e., $l_{1}=l_{2}=[\,]$) or a singleton set $[x]$ consisting of the first element $x$ of the queue, along with the updated queue. If $l_{1}=[\,]$, then the first element of the queue is computed by extracting the first element of the reverse of $l_{2}$. Given a list $l$ of elements as input, the main function enqueues all the elements of $l$, dequeues all the elements from the queue, and then asserts that the result equals $l$.
The correctness of the above program (i.e., the lack of assertion failures) is reduced to the satisfiability problem for the following CHCs.
$|x|=1\Rightarrow\mathit{Enq}(x,l_{1},l_{2},xl_{1},l_{2}).$
$\mathit{Deq}(\epsilon,\epsilon,\epsilon,\epsilon,\epsilon).$
$|x|=1\land\mathit{Reva}(l_{2},\epsilon,xl)\Rightarrow\mathit{Deq}(\epsilon,l_{2},x,l,\epsilon).$
$|x|=1\Rightarrow\mathit{Deq}(xl_{1}^{\prime},l_{2},x,l_{1}^{\prime},l_{2}).$
$\mathit{EnqAll}(\epsilon,l_{1},l_{2},l_{1},l_{2}).$
$l=xl^{\prime}\land|x|=1\land\mathit{Enq}(x,l_{1},l_{2},l_{1}^{\prime},l_{2}^{\prime})\land\mathit{EnqAll}(l^{\prime},l_{1}^{\prime},l_{2}^{\prime},r_{1},r_{2})\Rightarrow\mathit{EnqAll}(l,l_{1},l_{2},r_{1},r_{2}).$
$\mathit{Deq}(l_{1},l_{2},\epsilon,l_{1}^{\prime},l_{2}^{\prime})\Rightarrow\mathit{DeqAll}(l_{1},l_{2},\epsilon).$
$\mathit{Deq}(l_{1},l_{2},l,l_{1}^{\prime},l_{2}^{\prime})\land l\neq\epsilon\land\mathit{DeqAll}(l_{1}^{\prime},l_{2}^{\prime},r)\Rightarrow\mathit{DeqAll}(l_{1},l_{2},lr).$
$\mathit{EnqAll}(l,\epsilon,\epsilon,l_{1},l_{2})\land\mathit{DeqAll}(l_{1},l_{2},r)\Rightarrow r=l.$
Here, each predicate represents the relation between inputs and outputs of the corresponding function. For example, $\mathit{Enq}(x,l_{1},l_{2},r_{1},r_{2})$ means that $\mathtt{enq}\;x\;(l_{1},l_{2})$ may return $(r_{1},r_{2})$. The clauses for $\mathit{Reva}$ (which are found in Section 5.1) have been omitted.
The system of CHCs above has the following piecewise CSTP-model:
$\{\quad$
$\mathit{Reva}\mapsto(l_{1},l_{2},l_{1}^{R}l_{2}),$
$\mathit{Enq}\mapsto(l,l_{1},l_{2},l_{1},l^{R}l_{2}),$
$\mathit{Deq}\mapsto((\lambda(l_{1},l_{2},x,r_{1},r_{2}).l_{1}=l_{2}=\epsilon)\Rightarrow(\epsilon,\epsilon,\epsilon,\epsilon,\epsilon))$
$\qquad\ \land(\lambda(l_{1},l_{2},x,r_{1},r_{2}).l_{1}=\epsilon\land l_{2}\neq\epsilon)\Rightarrow(\epsilon,lx,x^{R},l^{R},\epsilon))$
$\qquad\ \land(\lambda(l_{1},l_{2},x,r_{1},r_{2}).l_{1}\neq\epsilon\land l_{2}\neq\epsilon)\Rightarrow(xl_{1},l_{2},x,l_{1},l_{2})),$
$\mathit{EnqAll}\mapsto(l,l_{1},l_{2},l_{1},l^{R}l_{2}),$
$\mathit{DeqAll}\mapsto(l_{1},l_{2},l_{1}l_{2}^{R})$
$\}.\quad$
Thus, if we set $\mathcal{P}$ to $\{p_{1},p_{2},p_{3},p_{4},p_{5}\}$ where:
$p_{1}=\lambda(l_{1},l_{2},l_{3},l_{4},l_{5}).l_{1}=l_{2}=\epsilon$
$p_{2}=\lambda(l_{1},l_{2},l_{3},l_{4},l_{5}).l_{1}=\epsilon\land l_{2}\neq\epsilon$
$p_{3}=\lambda(l_{1},l_{2},l_{3},l_{4},l_{5}).l_{1}\neq\epsilon\land l_{2}\neq\epsilon$
$p_{4}=\lambda(l_{1},l_{2},l_{3}).\mathit{true}$
$p_{5}=\lambda(l_{1},l_{2},l_{3},l_{4},l_{5}).\mathit{true},$
then we can automatically prove the satisfiability of the CHCs above. We expect that necessary predicates in $\mathcal{P}$ can typically be mined from the source program; in fact, in the above case, the function deq performs case analysis on whether $l_{1}$ and $l_{2}$ are empty lists. A general method to find an appropriate set $\mathcal{P}$ of predicates is left for future work.∎